# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ V Conclusion
In this work, we start with giving the set of operators that minimize the postselected symmetric error. We make an important observation about the error-minimizing measurements that such measurements never decide in favor of one of the possible states unless the prior probability has a particular value given by $(p_{\rho}^{*},p_{\sigma}^{*})$ in (14). For a lower prior probability of either of the states, any error-minimizing measurement never declares that state. This puts a fundamental question on minimizing postselected symmetric error $e(\Lambda)$ as a way to find the best measurement. After this, we have given an arbitrary construction of an error-minimizing measurement in terms of freely chosen variables. We then show by an example that the value of acceptance varies for different measurements taken from the set of error-minimizing measurements, despite all being error-minimizing. This example illustrates the need for maximizing the acceptance over the set of error-minimizing measurements. Further, we have given the expression of acceptance for an arbitrary error-minimizing measurement in terms of free variables used for constructing the error-minimizing measurement. This is followed by stating the expression for maximum acceptance in closed form-expression except for the case when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ by maximizing the expression of acceptance with respect to the free variable. Here, we would also add that, all these results are generalizable in a more simplified form in case of discriminating classical probability distributions on taking appropriate density matrices and measurements.
@@
The maximum acceptance obtained here over the set of error-minimizing measurements means that any lower value of acceptance is achievable with the same minimum error probability, but if higher acceptance is desired, the error probability is bound to increase. Thus, the work opens up the question of characterizing postselected symmetric error under the demand of higher acceptance. This work also opens up other new potential directions for research for example, studying acceptance in the asymptotic case. It may lead to a new class of quantum divergence parameterized by acceptance for any pair of quantum states. Recently postselected communication over quantum channels was proposed in[28] and capacity is derived. Some metric similar to acceptance in this work is potentially needed there as well. Similarly, methods of finding appropriate constrains can be generalized to problems on probabilistic protocols as in [29, 30, 31]. Overall, this work opens up a new arena to explore other matrices for performance of postselected symmetric hypothesis testing problems going beyond error probability and studying acceptance in related problems.
Appendix A Condition for equality
The appendix focuses on deriving certain certain relation between an operator $\nu\in\mathcal{P}(\mathcal{H})$ and projectors $\Pi_{\nu},\Pi_{\nu}^{\max}$, and $\Pi_{\nu}^{\min}$ and eigenvalues $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$. With the help of these properties we have given bounds on $\mathrm{Tr}\left(\zeta\nu\right)$ for $\zeta\in\mathcal{S}(\mathcal{H})$ or with more constraints on $\zeta$. These bounds are derived in Lemma A.3 and Lemma A.6, which will be utilized in deriving bounds in Appendix B in the proof of Theorem III.1.
We take $\nu\in\mathcal{P}(\mathcal{H})$ as fixed operator and $\zeta\in\mathcal{P}(\mathcal{H})$ as a variable operator with more constraints if needed. Eigenvalue decomposition of $\nu$ and projection operator of the subspace spanned by eigenvectors is given as
$\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|\text{ and }\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|,$ (37)
@@ Appendix A Condition for equality
where $k_{\nu,i}$ represents eigenvalue and $|e_{i}\rangle$ is corresponding eigenvector of $\nu$. The projection operators corresponding to the maximum and minimum eigenvalues are given as
$\Pi^{\max}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|\text{ and }\Pi^{\min}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\text{ respectively.}$. (38)
Recall that $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$ denote the maximum and minimum eigenvalue of $\nu$ respectively. We will use this notation throughout Appendix A.
@@
For any $\nu\in\mathcal{P}(\mathcal{H})$, the following holds:
  1. $(\mathrm{1})$
    $\Pi_{\nu}^{\max}\Pi_{\nu}=\Pi_{\nu}^{\max}$,
  2. $(\mathrm{2})$
    $\Pi_{\nu}^{\max}\nu=\|\nu\|_{\infty}\Pi_{\nu}^{\max}$,
  3. $(\mathrm{3})$
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=1\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$,
  4. $(\mathrm{4})$
    $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$,
  5. $(\mathrm{5})$
    $\Pi_{\nu}^{\min}\Pi_{\nu}=\Pi_{\nu}^{\min}$,
  6. $(\mathrm{6})$
    $\Pi_{\nu}^{\min}\nu=\|\nu\|_{\infty,0}\Pi_{\nu}^{\min}$,
  7. $(\mathrm{7})$
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=1\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$,
  8. $(\mathrm{8})$
    $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
@@ 8. $(\mathrm{8})$
Proof of each statement is given below.
  1. $(\mathrm{1})$
    Note that $\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}^{\max}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|$. Hence,
    $\Pi_{\nu}^{\max}\Pi_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}\sum_{j}|e_{i}\rangle\langle e_{i}||e_{j}\rangle\langle e_{j}|=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|=\Pi_{\nu}^{\max}.$
  2. $(\mathrm{2})$
    $\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}^{\max}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|$. Hence
    $\Pi_{\nu}^{\max}\nu$ =$=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}|e_{i}\rangle\langle e_{i}|k_{\nu,j}|e_{j}\rangle\langle e_{j}|=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}\mathds{1}(j=i)k_{\nu,j}|e_{j}\rangle\langle e_{j}|$
    =$=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}k_{\nu,i}|e_{i}\rangle\langle e_{i}|=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\|\nu\|_{\infty}|e_{i}\rangle\langle e_{i}|=\|\nu\|_{\infty}\Pi_{\nu}^{\max}.$
  3. $(\mathrm{3})$
    From the definition of set $\mathcal{S}(\Pi_{\nu}^{\max})$, it follows that $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\zeta$ and $\mathrm{Tr}\left(\zeta\right)=1$. So,
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\Pi_{\nu}\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\right)=\mathrm{Tr}\left(\zeta\right)=1.$ (39)
    First, third and fourth equality above follows from the definition of $\mathcal{S}(\Pi_{\nu}^{\max})$ and second equality follows from $(\mathrm{1})$.
  4. $(\mathrm{4})$
    Similar to the proof of $(\mathrm{3})$, using $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\zeta$ and $\mathrm{Tr}\left(\zeta\right)=1$, we get
    $\mathrm{Tr}\left(\zeta\nu\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\nu\right)=\|\nu\|_{\infty}\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\right)=\|\nu\|_{\infty}\mathrm{Tr}\left(\zeta\right)=\|\nu\|_{\infty}.$ (40)
    First, third and fourth equality above follows from the definition of $\mathcal{S}(\Pi_{\nu}^{\max})$ and second equality follows from $(\mathrm{2})$.
  5. $(\mathrm{5})$
    --$(\mathrm{8})$ Proof is similar to proof of $(\mathrm{1})-(\mathrm{4})$.
This completes the proof of all statements mentioned in the Lemma. ∎
Lemma A.2.
Given $\nu\in\mathcal{P}(\mathcal{H}),$$\langle\phi|\nu|\phi\rangle\leq\|\nu\|_{\infty}\ \forall\ |\phi\rangle\in\mathcal{H}$ and $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$.
Proof.
We first prove the inequality and then show the equivalence of equality. Using eigenvalue decomposition of $\nu$,
$\langle\phi|\nu|\phi\rangle=\sum_{i}\langle\phi|k_{\nu,i}|e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\langle\phi||e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}.$ (41)
Note that $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=\langle\phi|\left(\sum_{i}|e_{i}\rangle\langle e_{i}|\right)|\phi\rangle\leq\langle\phi|I|\phi\rangle=\langle\phi|\phi\rangle=1$ and $k_{\nu,i}\leq\|\nu\|_{\infty}\ \forall i$. Hence
$\langle\phi|\nu|\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\leq\sum_{i}\|\nu\|_{\infty}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty}\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\leq\|\nu\|_{\infty}.$ (42)
So $\langle\phi|\nu|\phi\rangle\leq\|\nu\|_{\infty}$, which is the inequality in the lemma.
For equality to hold, both the inequalities in (42) have to be equality. First inequality becomes equality iff $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$ and second inequality becomes equality iff $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1$. So we get
$\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1.$ (43)
Now, $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1\Leftrightarrow\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=1\Leftrightarrow\langle\phi|\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle=1\Leftrightarrow\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle=|\phi\rangle.$ So we get
$\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle.$ (44)
On substituting $\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$ in $|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle$, we get
$|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|\phi\rangle=\Pi_{\nu}^{\max}|\phi\rangle.$ (45)
For showing the equivalence, converse is also needed, which is proved below.
$|\phi\rangle=\Pi_{\nu}^{\max}|\phi\rangle=\!\!\!\!\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}\!\!\!\!|e_{i}\rangle\langle e_{i}|\phi\rangle\Rightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle.$ (46)
So, we get $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle\text{ with }\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$. Hence, $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$. ∎
Lemma A.3.
For any $\nu\in\mathcal{P}(\mathcal{H})$, $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}\ \forall\ \zeta\in\mathcal{D}(\mathcal{H})$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$.
Proof.
@@ 5. $(\mathrm{5})$
The proof is done in 2 steps. First we have shown that $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$, followed by showing $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}$. Combining these two, we get the stated result.
@@
    Take eigenvalue decomposition of $\zeta=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ with $\sum_{i}k_{\zeta,i}=1$. Then $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle$. Using Lemma A.2, $\langle\phi_{i}|\nu|\phi_{i}\rangle\leq\|\nu\|_{\infty}$ and $\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi_{i}\rangle=|\phi_{i}\rangle$. Hence,
@@
    Using statement $(\mathrm{4})$ of Lemma A.1, we get
    $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}.$ (49)
@@
$\zeta\in\mathcal{S}(\Pi)\Leftrightarrow\Pi\zeta=\zeta\Leftrightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$, where $\{|\phi_{i}\rangle\}$ is set of eigenvectors of $\zeta$.

### v2

@@ V Conclusions
In this work, we have given the set of all measurements that minimize the postselected symmetric error. We make an important observation about the error-minimizing measurements that such measurements never declare one of the possible states unless the prior probability has a particular value given by $(p_{\rho}^{*},p_{\sigma}^{*})$ in (16). For a lower prior probability of either of the states, any error-minimizing measurement never declares that state. So, if $p_{\rho}<p_{\rho}^{*}$, any error-minimizing measurement never declares the unknown state as $\rho$. This observation raises a fundamental question about minimizing the postselected symmetric error $e(\Lambda)$ as a means to search for the best measurement. Then, we have stated a method to construct an arbitrary error-minimizing measurement in terms of variables, which can be varied to obtain other error-minimizing measurements. Furthermore, we show by an example that the value of acceptance varies for different measurements taken from the set of error-minimizing measurements, despite all being error-minimizing. The example illustrates the need to maximize acceptance over the set of error-minimizing measurements. So, we have obtained the expression of acceptance for an arbitrary error-minimizing measurement in terms of free variables used to construct the error-minimizing measurement. Then, we have derived the maximum acceptance. We obtain the maximum acceptance for both sates in closed-form expression if prior probabilities $p_{\rho}\neq p_{\rho}^{*}$ and $p_{\sigma}\neq p_{\sigma}^{*}$, and as an optimization problem otherwise. These results are generalizable in a simplified form for discriminating classical probability distributions by taking appropriate density matrices and measurements.
@@
The maximum acceptance obtained in this work over the set of all error-minimizing measurements tells that any lower value of acceptance can be obtained while ensuring that the postselected symmetric error remains the same, i.e., achieves its minimum value. However, if a higher value of acceptance is desired, the postselected symmetric error will increase. So, this work opens up a problem of studying the trade-off between acceptance postselected symmetric error. Such a study has been done in the past with different metrics for error and acceptance [14], and requires an independent study in the context of the problem studied in our work. Such a study is likely to be useful in adaptive quantum hypothesis testing. This work also opens up other new directions for research, such as studying acceptance in the asymptotic case. It may lead to a new class of quantum divergence parameterized by acceptance. In the past, one-shot USD has been generalized to USD of unknown states, and separation increasing transformations [43, 7, 6]. PSD, studied in our work, can also be generalized in a similar fashion. [17] generalizes maximizing the success probability given an error margin for discrimination of states [18] to unitary operations. Work studied here can similarly be generalized to unitary operations and channels. Postselected communication over quantum channels was proposed in[27] and capacity is derived. A metric similar to acceptance in our work needs to be considered while obtaining capacity for postselected communication. The method of finding constraints on an arbitrary measurement for it to be an error-minimizing measurement, studied in our work, can be generalized to problems on probabilistic protocols, as in [35, 34, 32] to obtain the set of all operations that maximize the success probability. Overall, this work opens up a new arena to explore other matrices for the performance of postselected symmetric hypothesis testing problems, going beyond error probability and studying acceptance in related problems.
Appendix A Proof of Theorem 1
In this appendix, we take $\nu\in\mathcal{P}(\mathcal{H})$ as a fixed operator and $\zeta\in\mathcal{S}(\mathcal{H})$ as a variable operator. We find the maximum and the minimum value of $\mathrm{Tr}\left(\zeta\nu\right)$ by varying $\zeta\in\mathcal{S}(\mathcal{H})$ in Lemma A.1 and Lemma A.2 respectively. We also obtain the necessary and sufficient condition on $\zeta$ to obtain the maximum and the minimum value of $\mathrm{Tr}\left(\zeta\nu\right)$ in Lemma A.1 and Lemma A.2 respectively. We utilize these bounds, and condition on $\zeta$ for equality of these bounds in the proof of Theorem 1.
@@ Appendix A Proof of Theorem 1
Eigenvalue decomposition of $\nu$ and the projection operator of the subspace spanned by eigenvectors is given as $\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|$, where $k_{\nu,i}$ represents the eigenvalue and $|e_{i}\rangle$ is corresponding eigenvector of $\nu$. The projection operators the corresponding to the maximum and minimum eigenvalues are given as
$\Pi^{\max}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|\text{ and }\Pi^{\min}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\text{ respectively.}$ (36)
@@
For a given $\nu\in\mathcal{P}(\mathcal{H})$, we get $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}\ \forall\ \zeta\in\mathcal{D}(\mathcal{H})$ and
$\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max}).$
@@
The proof is done in 2 steps. First, we have shown that $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$, followed by showing $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}$.
@@
    Let us take the eigenvalue decomposition of $\zeta=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ with $\sum_{i}k_{\zeta,i}=1$. Hence, $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle$. Note that $\langle\phi_{i}|\nu|\phi_{i}\rangle\leq\|\nu\|_{\infty}$ from the definition of max norm. Further, $\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi_{i}\rangle=|\phi_{i}\rangle$ as for maximization of $\langle\phi_{i}|\nu|\phi_{i}\rangle$, $|\phi_{i}\rangle$ must lie in the eigenspace corresponding to the maximum eigenvalue of $\nu$. Hence,
@@
    Note that $\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|$, hence $\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}^{\max}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|$. Now, from the definition of set $\mathcal{S}(\Pi_{\nu}^{\max})$, it follows that $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\zeta$ and $\mathrm{Tr}\left(\zeta\right)=1$. So we get
    ${\zeta\nu}$ $={\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\nu}={\Pi_{\nu}^{\max}\zeta\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}|e_{i}\rangle\langle e_{i}|k_{\nu,j}|e_{j}\rangle\langle e_{j}|}$ (39)
    $={\Pi_{\nu}^{\max}\zeta\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}\mathds{1}(i=j)k_{\nu,j}|e_{j}\rangle\langle e_{j}|}$ (40)
    $={\Pi_{\nu}^{\max}\zeta\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}\|\nu\|_{\infty}|e_{j}\rangle\langle e_{j}|}=\|\nu\|_{\infty}{\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}}=\|\nu\|_{\infty}{\zeta}.$ (41)
    So we get $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow{\zeta\nu}=\|\nu\|_{\infty}{\zeta}\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\mathrm{Tr}\left(\zeta\right)=\|\nu\|_{\infty}$.
@@
For a given $\nu\in\mathcal{P}(\mathcal{H})$, we get $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu})$ and
$\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min}).$